OpenStudy (anonymous):

I would just like assistance with the problem, if possible. If a stone is tossed from the top of a 190 meter building, the height of the stone as a function of time is given by h(t)=-9.8t^2-10t +190, where t is in seconds, and height is in meters. After how many seconds will the stone hit the ground? Round to the nearest hundredth's place; include units in your answer. This is the help I had from two other people, but they ran off, lol. I do not know if it is finished or not. Help me please. So for the whole problem I have: h(t)=-9.8t^2 -10t +190 -9.8t^2-10t+190=0 t=-3.9 or t=3.9 D=b^2-4ac a

OpenStudy (saifoo.khan):

i solved that for u.

OpenStudy (anonymous):

hey saifoo

OpenStudy (saifoo.khan):

i even remember the method, lol.

OpenStudy (saifoo.khan):

Hi Maggie.

OpenStudy (anonymous):

show me! i dunno how to solve

OpenStudy (anonymous):

so u got dissed by polpak to huh

OpenStudy (bahrom7893):

Again? Didn't we do this like 4 times?

OpenStudy (anonymous):

yup;(

OpenStudy (anonymous):

its ok ur not the only 1

OpenStudy (saifoo.khan):

Maggie, lol. v wil talk about it sometime later. If i talk now, my answers will be deleted.

OpenStudy (anonymous):

yeah, but everyone left, what is the final answer? isi t just t=-3.9 or t=3.9? or is it d=7548?

OpenStudy (anonymous):

k

OpenStudy (saifoo.khan):

WRONG.

OpenStudy (saifoo.khan):

answers were: t = 1.02 or t = 0.

OpenStudy (bahrom7893):

t=3.9

OpenStudy (anonymous):

how u get that answer?

OpenStudy (bahrom7893):

SAIFOO YOU ARE WRONG the answer is t = 3.9 u already said it.. set the equation equal to 0

OpenStudy (saifoo.khan):

but the height is given.

OpenStudy (saifoo.khan):

190 will be cancelled with 190.

OpenStudy (anonymous):

the other one said you missed a part of it. this part. -9.8t^2 -10t +190=0

OpenStudy (bahrom7893):

h(t)=-9.8t^2 -10t +190 h(t) = 0, when the ball touches the ground so -9.8t62 - 10t + 190 = 0

OpenStudy (saifoo.khan):

190 =-9.8t^2 -10t +190 190 - 190=-9.8t^2 -10t

OpenStudy (anonymous):

bahrom where did the 62 come from?

OpenStudy (saifoo.khan):

-9.8t^2 -10t

OpenStudy (bahrom7893):

t^2 sorry i forgot to press shift when doing ^

OpenStudy (anonymous):

so that is the final answer, this is what I have so far, I will need to come back and give you the answer, just a minute. stay here.

OpenStudy (bahrom7893):

-9.8t^2 - 10t + 190 = 0, you could've guessed that D = b^2 - 4ac = 100 + 4*9.8*190 = 100 + 7448 = 7548

OpenStudy (saifoo.khan):

i didnt used Determinate. :p

OpenStudy (bahrom7893):

t1 = (10 - sqrt(7548))/(2(-9.8)) t2 = (10 + sqrt(7548))/(2(-9.8)) Saifoo first of all it's called discriminant, second of all your method is wrong.

OpenStudy (anonymous):

once u settle ur debate, could one of u xplain how u do the problems?

OpenStudy (bahrom7893):

t1 = 3.92240952 t2 = -4.94281768 We disregard the negative sign.

OpenStudy (bahrom7893):

Gandalfwiz aren't u reading my posts?

OpenStudy (anonymous):

So for the whole problem I have: h(t)=-9.8t^2 -10t +190 -9.8t^2-10t+190=0 t=-3.9 or t=3.9 D=b^2-4ac a=-9.8 b=-10 c= 190 -10^2-4*-9.8*190 100-4*-9.8 100+39.2*190 100+7448 =7548 Is this right? So what is the final answer t=-3.9 or t=3.9? Or is it 7548?

OpenStudy (bahrom7893):

We disregrad the time with the negative sign since time cannot be negative, therefore the answer is: t1 = 3.92240952

OpenStudy (anonymous):

sorry ! I'm Back! I'm back!

OpenStudy (bahrom7893):

For the 7th and last time, here is the complete, step-by-step solution: h(t)=-9.8t^2 -10t +190 h(t) = 0, when the ball touches the ground so -9.8t^2 - 10t + 190 = 0 D = b^2 - 4ac = 100 + 4*9.8*190 = 100 + 7448 = 7548 t1 = (10 - sqrt(7548))/(2(-9.8)) t2 = -4.94281768 We disregrad the time with the negative sign since time cannot be negative, therefore the answer is: t1 = 3.92240952

OpenStudy (bahrom7893):

Oh to the nearest hundredth that is: t = 3.92

OpenStudy (anonymous):

How did you come up with your t1 and t2 answers?

OpenStudy (anonymous):

Ok, I got it Thank you for your help. See the other ones I posted, I never got the whole problems, just pieces of it.

OpenStudy (bahrom7893):

Nickie do u know how to do quadratics? ax^2+bx+c=0 D = b^2-4ac x1 = (-b-sqrt(D))/(2a) x2 = (-b+sqrt(D))/(2a)

OpenStudy (bahrom7893):

finally..

OpenStudy (anonymous):

I was copying the problem. you added the 4 in the diterminate formula. the formula says D=b^2-4ac doesnt that make it all wrong. Nevermind you did it right, just not written down right. Thank you for your help bahrom.